OpenStudy (mirandaregina):

Write an equation of the line that passes through the given point and satisfies the given condition. (-3, -5); parallel to y = -4x + 1

OpenStudy (anonymous):

eq. of line parallel to y=-4x+1 is y=-4x+c it passes through (-3,-5) -5=-4*-3+c c=-5-12=-17 reqd. eq. is y=-4x-17

OpenStudy (mirandaregina):

That's the answer I got as well. Thank you, surjithayer.

OpenStudy (anonymous):

yw.